TIFFDUMP(1) TIFFDUMP(1) NAME tiffdump−printverbatiminformationaboutTIFFfiles SYNOPSIS tiffdump[options]name... DESCRIPTION tiffdumpdisplaysdirectoryinformationfromfilescreatedaccordingtotheTagImageFileFormat, Revision6.0.TheheaderofeachTIFFfile(magicnumber,version,andfirstdirectoryoffset)is displayed,followedbythetagcontentsofeachdirectoryinthefile.Foreachtag,thename, datatype,count,andvalue(s)isdisplayed.Whenthesymbolicnameforatagordatatypeis known,thesymbolicnameisdisplayedfollowedbyit’snumeric(decimal)value.Tagvaluesare displayedenclosedin‘‘<>’’charactersimmediatelyprecededbythevalueofthecountfield.For example,anImageWidthtagmightbedisplayedas‘‘ImageWidth(256)SHORT(3)1<800>’’. tiffdumpisparticularlyusefulforinvestigatingthecontentsofTIFFfilesthatlibtiffdoesnot understand. OPTIONS −hForcenumericdatatobeprintedinhexadecimalratherthanthedefaultdecimal. −oDumpthecontentsoftheIFDattheaparticularfileoffset.ThefileoffsetmaybespecifiedusingtheusualC-stylesyntax;i.e.aleading‘‘0x’’forhexadecimalandaleading‘‘0’’ foroctal. SEEALSO tiffinfo(1),libtiff(3) October15,1995 1